# Efficient detection of entanglement by stimulated disentanglement

### v1

@@
Standard detection of entanglement relies on local measurements of the individual particles, evaluating their correlations in post-processing. For time-energy entangled photons, either times $(t_{1},t_{2})$ or energies $(E_{1},E_{2})$ are measured, but not both due to the mutual quantum uncertainty, providing only partial information of the entanglement. In principle, a global detector could recover the complete information of entanglement in a single shot if it could measure the combined correlated variables $(t_{1}-t_{2})$ and $(E_{1}+E_{2})$ without measuring the individual energies or times. Such a global measurement is possible using the reverse disentangling interaction, like sum-frequency generation (SFG), but nonlinear interactions at the single-photon level are ridiculously inefficient. We overcome this barrier by stimulating the nonlinear SFG interaction with a strong pump, thereby measuring both the energy-sum (SFG spectrum) and the time-difference (response to group delay/dispersion) simultaneously and efficiently. We generate bi-photons with extreme time-energy entanglement (octave-spanning spectrum of 113THz) and measure a relative uncertainty between time-difference and energy-sum of $\Delta(t_{1}-t_{2})\Delta(E_{1}+E_{2})\approx\!2\!\cdot\!10^{-13}h$, , violating the classical bound by >12 orders of magnitude. The presented coherent SFG dramatically enhances the detection SNR compared to standard methods since it ideally rejects erroneous coincidences in both time and energy, paving the way for sensing applications, such as quantum illumination (radar) and more.
@@ 1 Introduction
The generation of entanglement relies on an entangling interaction between the parties, which for photons inherently entails a nonlinearity, such as three- or four-wave mixing. A well-known example is the generation of entangled photons by spontaneous parametric down-conversion (SPDC) in a three-waves mixing medium, where a pump ($p$) photon from an intense laser splits into a pair of lower-energy photons, named signal ($s$) and idler ($i$) [13, 14]. These generated bi-photons inherit the properties of the pump, indicating that the sum of all their preserved properties, like energies (frequencies), linear momenta ($k$-vectors), angular momenta (transverse modes) or spins (polarizations) equal those of the original pump, but the properties of each of the photons are not defined and can span a range of possible values. In this paper, we focus on time-energy entangled pairs of photons that are generated by broadband spontaneous parametric down-conversion from a narrowband pump laser: The total energy is fixed ($\nu_{s}+\nu_{i}=\nu_{p}$), as dictated by the pump, but the individual photon energies vary across a very wide bandwidth, limited only by the phase-matching of the nonlinear medium, which can span up to an octave of bandwidth ($\Delta\nu\!\approx\!113$THz in our experiment). Conversely, the nonlinear generation of the pair is “instantaneous”, indicating that the arrival times of the photons are tightly correlated down to the Fourier limit $t_{s}\!-\!t_{i}\!\leq\!1\!/\!\Delta\nu$ ($\approx\!9$ fs in our experiment).
@@
whereas quantum mechanically no fundamental limit exists. Indeed, for narrowband bi-photons, where the time and frequency correlations could be measured directly with standard detectors, $\Delta(\nu_{1}\!+\!\nu_{2})\Delta(t_{1}\!-\!t_{2})\!=\!0.063$ was previously observed [15], violating this bound by a factor of $\sim\!8$. In comparison, broadband bi-photons that are generated from a narrowband (single-frequency) pump can dramatically violate the bound of Eq. 1, since their time-correlation can be extremely short ($\sim\!9$fs in our experiment), whereas the observed correlation in frequency will be limited only by the spectral resolution of the sum-frequency measurement. In our experiment, reported hereon, we demonstrate violation of Eq. 1 by nearly $13$ orders of magnitude (!).
@@
Since efficient global detection was not available so far, the standard characterization of entanglement relies on correlations of local measurements. In discrete-variable systems (such as polarization-entangled photons), Bell’s inequalities utilize measurements in carefully selected bases to quantify the quantum violation of the classical correlation bounds [16, 17]. However, the direct extension of Bell’s inequalities to continuous-variable (CV) systems remains an ongoing challenge, both theoretically and experimentally. Many important techniques were proposed, including continuous variables (CV) Bell inequalities, e.g. [18, 19, 20, 21, 22], phase-space reconstructions [23, 24], variance-based inequalities (i.e. the Simon-Duan criterion and its generalization to higher moments [25, 26, 27, 28, 29], and more general Gaussian-state criteria [30], but no single method can universally and conclusively detect all forms of (CV) entanglement. Moreover, some of the above methods are challenging in practice, involving complex preparation of sophisticated measurements, and may result in very small Bell violations [31]. Moreover, limitations introduced by detector inefficiencies, optical losses, or finite sampling rates can obscure genuine quantum correlations. Nevertheless, various techniques were developed to witness and quantify time-energy entanglement, as summarized in Fig. 1, including coincidence detection in time with Hong-Ou-Mandel (HOM) interference [32], spectral-correlation measurement using SU(1,1) interferometry [33, 34] and direct SFG [35, 36, 37]. While these methods have advanced the field, none of them provided the complete correlation efficiently from both time and energy.
@@
Figure 1: (a) Detection methods of time-energy entanglement: Coincidence detection with HOM interference, Direct SFG and SU(1,1) Interference are shown in the top three rows, in comparison to the new Coherent SFG (fourth row) that we present now. In all schemes, a pump laser (black line) interacts with a nonlinear medium ($\chi^{(2)}$) to generate entangled photon pairs—signal (blue) and idler (red), followed by manipulation and measurement. The manipulation can includes removal of the pump or phase modulation (or both). Table (b) on the right highlights the pros and cons of each method regarding four critical aspects: access to the time correlation, access to the energy correlation, detection efficiency, and whether the detectable flux of entangled photons is limited by the speed of the photo-detectors.
@@
Conceptually, a global measurement of time-difference and energy-sum is achievable via time-reversal of the nonlinear entangling interaction. Consider sum frequency generation (SFG - see Fig. 1b) - the inverse of parametric down conversion (PDC) [38], where a pair of signal and idler photons is annihilated to form a single SFG photon, to be later detected. Measuring the energy of the SFG photon provides only the energy-sum of the entangled pair without the individual energies. Conversely, SFG requires both photons to arrive simultaneously, thereby acting as a coincidence detector of $t_{s}\!-\!t_{i}\!\approx\!0$ which is blind to the specific times. Compared to standard methods of coincidence, HOM or SU(1,1) interference, an efficient SFG detector will dramatically enhance the sensitivity and SNR, since it will “doubly reject” accidental coincidences in both energy and time.
@@
Furthermore, the enhanced SNR of SFG detection directly indicates the time-energy entanglement: While the bi-photon amplitude of each individual signal-idler frequency-pair in the spectrum is well below the detection threshold of $1\sqrt{\text{photons}}/\text{mode}$, the coherent sum of all modes across the spectrum can generate a well-detectable SFG signal with practically no noise background due to the ideal rejection of accidental coincidences (as we discuss hereon). Consequently, SFG is recognized as ideal for quantum sensing applications that rely on entanglement, such as quantum radar and imaging with quantum illumination [39, 37]. Unfortunately, the efficiency of direct SFG at the single-photon level is dramatically low (typically $10^{-8}$ or lower). Thus, although SFG of entangled photons was demonstrated to expose the non-classical features of time-energy entangled photons [36]), it cannot form a practical detector for real-world applications. Table 1 summarizes the advantages and limitations of the existing methods of entanglement detection.
@@
To overcome the inefficiency barrier, we propose to stimulate the nonlinear interaction in the detection crystal with a strong coherent pump field, which dramatically enhances the SFG efficiency of photon-pairs to near unity [40]. Coherent stimulation—now phase-dependent—underpins nonlinear SU(1,1) interference, where the relative phase between the pump field and bi-photons determines the energy flow in the second nonlinear crystal (from bi-photons to pump or vice versa). This non-classical SU(1,1) interference is also used for bi-photon detection (Fig. 1c) by measuring the spectral intensity of bi-photons after the detection crystal. The resulting high-contrast interference fringes reveal the nonlinear pairwise interference of photons, with fringe contrast serving as a non-classical witness [33]. Remarkably, this process remains efficient even in lossy conditions, selectively extracting correlated bi-photons. Moreover, SU(1,1) coherence enabled parametric homodyne detection [41], which uses the pump as a common phase reference to measure a specified two-mode quadrature of all signal-idler frequency-pairs across a broad spectrum. And yet, SU(1,1) detection does not capture the complete quantum correlation, since it relies on local measurements of the down-converted photons.
@@
To obtain full correlation information, we return to SFG detection with a twist: coherent SFG (Fig. 1d). This method leverages the near unit-efficiency of SU(1,1) interference while shifting detection from the bi-photons back to the pump—measuring the pump’s depletion, gain, or phase shift from the coherent SFG conversion. In doing so, it combines SFG’s global, noise-rejecting nature with the efficiency of SU(1,1) interference.
@@ 2 Experiment and Results
Aiming to demonstrate a highly efficient sensor of correlation for both time and energy simultaneously, we are confronted with three major challenges: First, rejection of the high intensity pump that otherwise would overwhelm the detector; Second, precise control of the spectral phase and material dispersion to ensure that the entire spectrum of entangled bi-photons contributes constructively to the measured SFG signal; and third, stabilization of the relative phase between the SFG and the pump for a reliable coherent measurement.
@@
A simplified layout of our experimental configuration is shown in Fig. 2, illustrating the key components and their roles. The configuration is divided into five functional blocks: The laser sources (green) that include the pump ($\lambda$=885 nm) for SPDC and the seed ($\lambda$=1560 nm) that is added for stabilization of the SPDC phase relative to the pump (as detailed below); The nonlinear crystal for PDC and SFG, placed within a Sagnac interferometer (brown) for efficient rejection of the large pump background; This allows the SFG detection (yellow) with lock-in measurement to isolate the SFG contribution only; The diagnostics and phase lock (purple) measure the PDC spectrum at the output of the SU(1,1) interferometer and extracts the phase of the nonlinear interaction in real time for stabilization; And finally, control of the spectral phase of the bi-photons (blue) that include a prism-based pulse shaper for precise tuning of the group delay / dispersion between the photons, as well as a fast piezo-actuated mirror for modulation and stabilization of the PDC / SFG phase relative to the pump to allow the lock-in detection of the SFG. The full detection methodology is detailed in the Methods below.
@@
The main results are shown in Fig. 3, presenting the correlation in both energy-sum and time-difference (or group delay). The top panel (a) reflects the SFG spectrum of recombination, highlighting the precise match of the sum-energy of the bi-photons with the original pump energy and confirming that the process is dominated by the recombination of correlated photons back to the pump frequency $\nu_{SFG}\!=\!\nu_{p}$, while the probability of accidental SFG to other SFG frequencies from non-twin photons vanishes (as theoretically predicted, see below). Specifically, panel (a) shows the RF spectrum of the of the detected pump+SFG intensity, demonstrating a single peak of high contrast ($>\!37$dB) at the lock-in dither frequency $\omega$. The measured intensity modulation is due to the beat of interference between the carrier pump at $\nu_{P}$ and the SFG sidebands at $\nu_{SFG}\!=\!\nu_{P}\!\pm\!\omega$ due to the phase modulation of the SPDC, thereby capturing the frequency-difference between the SFG signal and the pump at Hz-level resolution ($20$Hz width), limited only by the integration time of the RF analyzer. The background of the SFG peak is purely the electronic noise floor of our digital spectrum analyzer (Moku:Lab). The absence of an optical contribution to the background at other frequencies $\nu\!\neq\!\omega$ indicates that the recombination of correlated bi-photons dominates and that accidental SFG events are insignificant.
@@
The central panel (b) of Fig. 3 reflects the correlation in time-difference between the photons by observing the dependence of the SFG spectral peak (at the dither frequency $\nu$) on the relative group delay between the signal and the idler photons before entering the SFG crystal. We continuously tune the group delay (and GDD) by varying the amount of material that the beam of bi-photons traverses in the spectral-shaper (by shifting the Sapphire prisms in/out). The additional material dispersion imprints a quadratic phase $\phi(\nu)\!=\!\beta_{2}\nu^{2}$ on the broad spectrum of spontaneous bi-photons, which in time indicates a group delay (and dispersion) between each signal frequency and its twin idler.; and the SFG signal extracted from the lock-in detection is simply $I_{SFG}\!\sim\!\text{Re}[\alpha_{r}^{*}\sum_{\nu}a_{s}(\nu)a_{i}(-\nu)\mathrm{e}^{\mathrm{i}\phi(\nu)}]$, which reflects a homodyne measurement of the SFG sum over all the signal-idler pairs with $\alpha_{r}$ the coherent amplitude of the pump residue at the Sagnac output (see derivation below). Since even small GDD values alter the phase of the spontaneous SFG field, the variation of the GDD is immediately reflected in the homodyne, causing the sharp peak. The bottom panels (c) and (d) show the spectral SU(1,1) interferograms for two specific GDD values, marked A and B on the central panel (b): A - near GDD compensation (bottom right, blue), where the spectral phase is flat; and B - at a large GDD value (bottom left, violet), where the large quadratic spectral phase is imprinted on the spectral fringes.
@@
Finally, all the GDD-scans of Fig. 4 with various residue phases demonstrate a sharp peak near the zero-GDD point, which is a direct signature of the time correlation of the photon pairs. Specifically, the peak width $\delta\beta\!\approx\!4\pi/\Delta\nu^{2}\approx 30\,\mathrm{fs}^{2}$ is inversely proportional to the bi-photons bandwidth of $\Delta\nu\!=\!113$ THz squared (for $\varphi_{r}\!=\!0,\pi$ the peak is split). From a physical point of view, this width is also the GDD tolerance for the photon pairs to keep coherence, which reflects their time correlation of the bi-photons at $\Delta t\!=\!1/\Delta\nu\!\approx\!9$ fs. Together with the measured linewidth of $20$Hz for the SFG peak, mentioned above we obtain $\Delta(\nu_{1}\!+\!\nu_{2})\Delta(t_{1}\!-\!t_{2})\!\approx\!2\!\cdot\!10^{-13}$, showing a giant violation of the classical bound (Eq. 1). The longer-range oscillations in the GDD-scans beyond the central coherent SFG peak are mainly due to the fact that we scan the GDD, rather than the group delay, which indicates that the spectral phase at the center of the spectrum varies quadratically and not linearly across the spectrum.

### v2

@@
Standard detection of entanglement relies on local measurements of the individual particles, evaluating their correlations in post-processing. For time-energy entangled photons, either times $(t_{1},t_{2})$ or energies $(E_{1},E_{2})$ are measured, but not both due to the mutual quantum uncertainty, providing only partial information of the entanglement. In principle, a global detector could recover the complete information of entanglement in a single shot if it could measure the combined correlated variables $(t_{1}-t_{2})$ and $(E_{1}+E_{2})$ without measuring the individual energies or times. Such a global measurement is possible using the reverse disentangling interaction, like sum-frequency generation (SFG), but nonlinear interactions at the single-photon level are exceedingly inefficient. We overcome this barrier by stimulating the nonlinear SFG interaction with a strong pump, thereby measuring both the energy-sum (SFG spectrum) and the time-difference (response to group delay/dispersion) simultaneously and efficiently. We generate biphotons with extreme time-energy entanglement (octave-spanning spectrum of 113THz) and measure a relative uncertainty between time-difference and energy-sum of $\Delta(t_{1}-t_{2})\Delta(E_{1}+E_{2})\approx\!2\!\cdot\!10^{-13}h$, violating the classical bound by more than 12 orders of magnitude. The presented coherent SFG dramatically enhances the detection SNR compared to standard methods since it ideally rejects erroneous coincidences in both time and energy, paving the way for sensing applications, such as quantum illumination (radar) and more.
@@ 1 Introduction
The generation of entanglement relies on an entangling interaction between the parties, which for photons inherently entails a nonlinearity, such as three- or four-wave mixing. A well-known example is the generation of entangled photons by spontaneous parametric down-conversion (SPDC) in a three-wave mixing medium, where a pump ($p$) photon from an intense laser splits into a pair of lower-energy photons, named signal ($s$) and idler ($i$) [13, 14]. These generated biphotons inherit the properties of the pump, indicating that the sum of all their preserved properties, like energies (frequencies), linear momenta ($k$-vectors), angular momenta (transverse modes) or spins (polarizations) equal those of the original pump, but the properties of each of the photons are not defined and can span a range of possible values. In this paper, we focus on time-energy entangled pairs of photons that are generated by broadband spontaneous parametric down-conversion from a narrowband pump laser: The total energy is fixed ($\nu_{s}+\nu_{i}=\nu_{p}$), as dictated by the pump, but the individual photon energies vary across a very wide bandwidth, limited only by the phase-matching of the nonlinear medium, which can span up to an octave of bandwidth ($\Delta\nu\!\approx\!113$THz in our experiment). Conversely, the nonlinear generation of the pair is “instantaneous”, indicating that the arrival times of the photons are tightly correlated down to the Fourier limit $t_{s}\!-\!t_{i}\!\leq\!1\!/\!\Delta\nu$ ($\approx\!9$ fs in our experiment).
@@
whereas quantum mechanically no fundamental limit exists. Indeed, for narrowband biphotons, where the time and frequency correlations could be measured directly with standard detectors, $\Delta(\nu_{1}\!+\!\nu_{2})\Delta(t_{1}\!-\!t_{2})\!=\!0.063$ was previously observed [15], violating this bound by a factor of $\sim\!8$. In comparison, broadband biphotons that are generated from a narrowband (single-frequency) pump can dramatically violate the bound of Eq. 1, since their time-correlation can be extremely short ($\sim\!9$fs in our experiment), whereas the observed correlation in frequency will be limited only by the spectral resolution of the sum-frequency measurement. In our experiment, reported hereon, we demonstrate violation of Eq. 1 by nearly $13$ orders of magnitude (!).
@@
Since efficient global detection was not available so far, the standard characterization of entanglement relies on correlations of local measurements. In discrete-variable systems (such as polarization-entangled photons), Bell’s inequalities utilize measurements in carefully selected bases to quantify the quantum violation of the classical correlation bounds [16, 17]. However, the direct extension of Bell’s inequalities to continuous-variable (CV) systems remains an ongoing challenge, both theoretically and experimentally. Many important techniques were proposed, including continuous variables (CV) Bell inequalities, e.g. [18, 19, 20, 21, 22], phase-space reconstructions [23, 24], variance-based inequalities (i.e. the Simon-Duan criterion and its generalization to higher moments [25, 26, 27, 28, 29], and more general Gaussian-state criteria [30], but no single method can universally and conclusively detect all forms of (CV) entanglement. Moreover, some of the above methods are challenging in practice, involving complex preparation of sophisticated measurements, and may result in very small Bell violations [31]. Moreover, limitations introduced by detector inefficiencies, optical losses, or finite sampling rates can obscure genuine quantum correlations. Nevertheless, various techniques were developed to witness and quantify time-energy entanglement, as summarized in Fig. 1a, including coincidence detection in time with Hong-Ou-Mandel (HOM) interference [32], spectral-correlation measurement using SU(1,1) interferometry [33, 34] and direct sum-frequency generation (SFG) [35, 36, 37]. While these methods have advanced the field, none of them provided the complete correlation efficiently from both time and energy.
@@
Figure 1: (a) Detection methods of time-energy entanglement: Coincidence detection with HOM interference, Direct SFG and SU(1,1) Interference are shown in the top three rows, in comparison to the new Coherent SFG (fourth row) that we present now. In all schemes, a pump laser (black line) interacts with a nonlinear medium ($\chi^{(2)}$) to generate entangled photon pairs—signal (blue) and idler (red), followed by manipulation and measurement. The manipulation can includes removal of the pump or phase modulation (or both). (b) Pros and cons of each method: Four critical aspects are addressed – access to the time correlation $t_{1}-t_{2}$, access to the energy correlation $\omega_{1}+\omega_{2}$, detection efficiency, and whether the detectable flux of entangled photons is limited by the speed of the photo-detectors.
@@
Conceptually, a global measurement of time-difference and energy-sum is achievable via time-reversal of the nonlinear entangling interaction. Consider SFG (Fig. 1a) - the inverse of parametric down conversion (PDC) [38], where a pair of signal and idler photons is annihilated to form a single SFG photon, to be later detected. Measuring the energy of the SFG photon provides only the energy-sum of the entangled pair without the individual energies. Conversely, SFG requires both photons to arrive simultaneously, thereby acting as a coincidence detector of $t_{s}\!-\!t_{i}\!\approx\!0$ which is blind to the specific times. Compared to standard methods of coincidence, HOM or SU(1,1) interference, an efficient SFG detector will dramatically enhance the sensitivity and SNR, since it will “doubly reject” accidental coincidences in both energy and time.
@@
Furthermore, the enhanced SNR of SFG detection directly indicates the time-energy entanglement: While the biphoton amplitude of each individual signal-idler frequency-pair in the spectrum is well below the detection threshold of $1\sqrt{\text{photons}}/\text{mode}$, the coherent sum of all modes across the spectrum can generate a well-detectable SFG signal with practically no noise background due to the ideal rejection of accidental coincidences (as we discuss hereon). Consequently, SFG is recognized as ideal for quantum sensing applications that rely on entanglement, such as quantum radar and imaging with quantum illumination [39, 37]. Unfortunately, the efficiency of direct SFG at the single-photon level is dramatically low (typically $10^{-8}$ or lower). Thus, although SFG of entangled photons was demonstrated to expose the non-classical features of time-energy entangled photons [36]), it cannot form a practical detector for real-world applications. The table in Fig. 1b summarizes the advantages and limitations of the existing methods of entanglement detection.
@@
To overcome the inefficiency barrier, we propose to stimulate the nonlinear interaction in the detection crystal with a strong coherent pump field, which dramatically enhances the SFG efficiency of photon-pairs to near unity [40]. Coherent stimulation—now phase-dependent—underpins nonlinear SU(1,1) interference, where the relative phase between the pump field and biphotons determines the energy flow in the second nonlinear crystal (from biphotons to pump or vice versa). This non-classical SU(1,1) interference is also used for biphoton detection (Fig. 1a) by measuring the spectral intensity of biphotons after the detection crystal. The resulting high-contrast interference fringes reveal the nonlinear pairwise interference of photons, with fringe contrast serving as a non-classical witness [33]. Remarkably, this process remains efficient even in lossy conditions, selectively extracting correlated biphotons. Moreover, SU(1,1) coherence enabled parametric homodyne detection [41], which uses the pump as a common phase reference to measure a specified two-mode quadrature of all signal-idler frequency-pairs across a broad spectrum. And yet, SU(1,1) detection does not capture the complete quantum correlation, since it relies on local measurements of the down-converted photons.
@@
To obtain full correlation information, we return to SFG detection with a twist: coherent SFG (Fig. 1a). This method leverages the near unit-efficiency of SU(1,1) interference while shifting detection from the biphotons back to the pump—measuring the pump’s depletion, gain, or phase shift from the coherent SFG conversion. In doing so, it combines SFG’s global, noise-rejecting nature with the efficiency of SU(1,1) interference.
@@ 2 Experiment and Results
Aiming to demonstrate a highly efficient sensor of correlation for both time and energy simultaneously, we are confronted with three major challenges: First, rejection of the high intensity pump that otherwise would overwhelm the detector; Second, precise control of the spectral phase and material dispersion to ensure that the entire spectrum of entangled biphotons contributes constructively to the measured SFG signal; and third, stabilization of the relative phase between the SFG and the pump for a reliable coherent measurement.
@@
A simplified layout of our experimental configuration is shown in Fig. 2, illustrating the key components and their roles. The configuration is divided into five functional blocks: The laser sources (green) that include the pump ($\lambda$=885 nm) for SPDC and the seed ($\lambda$=1560 nm) that is added for stabilization of the SPDC phase relative to the pump (as detailed below); The nonlinear crystal for PDC and SFG, placed within a Sagnac interferometer (brown) for efficient rejection of the large pump background; This allows the SFG detection (yellow) with lock-in measurement to isolate the SFG contribution only; The diagnostics and phase lock (purple) measure the PDC spectrum at the output of the SU(1,1) interferometer and extracts the phase of the nonlinear interaction in real time for stabilization; And finally, control of the spectral phase of the biphotons (blue) that include a prism-based pulse shaper for precise tuning of the group delay / dispersion between the photons, as well as a fast piezo-actuated mirror for modulation and stabilization of the PDC / SFG phase relative to the pump to allow the lock-in detection of the SFG. The full detection methodology is detailed in the Methods below.
@@
The main results are shown in Fig. 3, presenting the correlation in both energy-sum and time-difference (or group delay). The top panel (a) reflects the SFG spectrum of recombination, highlighting the precise match of the sum-energy of the biphotons with the original pump energy and confirming that the process is dominated by the recombination of correlated photons back to the pump frequency $\nu_{SFG}\!=\!\nu_{p}$, while the probability of accidental SFG to other SFG frequencies from non-twin photons vanishes (as theoretically predicted, see below). Specifically, panel (a) shows the RF spectrum of the of the detected pump+SFG intensity, demonstrating a single peak of high contrast ($>\!37$dB) at the lock-in dither frequency $\omega$. The measured intensity modulation is due to the beat of interference between the carrier pump at $\nu_{P}$ and the SFG sidebands at $\nu_{SFG}\!=\!\nu_{P}\!\pm\!\omega$ due to the phase modulation of the SPDC, thereby capturing the frequency-difference between the SFG signal and the pump at Hz-level resolution ($20$Hz width), limited only by the integration time of the RF analyzer. The background of the SFG peak is purely the electronic noise floor of our digital spectrum analyzer (Moku:Lab). The absence of an optical contribution to the background at other frequencies $\nu\!\neq\!\omega$ indicates that the recombination of correlated biphotons dominates and that accidental SFG events are insignificant.
@@
The central panel (b) of Fig. 3 reflects the correlation in time-difference between the photons by observing the dependence of the SFG spectral peak (at the dither frequency $\nu$) on the relative group delay between the signal and the idler photons before entering the SFG crystal. We continuously tune the group delay (and GDD) by varying the amount of material that the beam of biphotons traverses in the spectral-shaper (by shifting the Sapphire prisms in/out). The additional material dispersion imprints a quadratic phase $\phi(\nu)\!=\!\beta_{2}\nu^{2}$ on the broad spectrum of spontaneous biphotons, which in time indicates a group delay (and dispersion) between each signal frequency and its twin idler.; and the SFG signal extracted from the lock-in detection is simply $I_{SFG}\!\sim\!\text{Re}[\alpha_{r}^{*}\sum_{\nu}a_{s}(\nu)a_{i}(-\nu)\mathrm{e}^{\mathrm{i}\phi(\nu)}]$, which reflects a homodyne measurement of the SFG sum over all the signal-idler pairs with $\alpha_{r}$ the coherent amplitude of the pump residue at the Sagnac output (see derivation below). Since even small GDD values alter the phase of the spontaneous SFG field, the variation of the GDD is immediately reflected in the homodyne, causing the sharp peak. The bottom panels (c) and (d) show the spectral SU(1,1) interferograms for two specific GDD values, marked A and B on the central panel (b): A - near GDD compensation (bottom right, blue), where the spectral phase is flat; and B - at a large GDD value (bottom left, violet), where the large quadratic spectral phase is imprinted on the spectral fringes.
@@
Finally, all the GDD-scans of Fig. 4 with various residue phases demonstrate a sharp peak near the zero-GDD point, which is a direct signature of the time correlation of the photon pairs. Specifically, the peak width $\delta\beta\!\approx\!4\pi/\Delta\nu^{2}\approx 30\,\mathrm{fs}^{2}$ is inversely proportional to the biphotons bandwidth of $\Delta\nu\!=\!113$ THz squared (for $\varphi_{r}\!=\!0,\pi$ the peak is split). From a physical point of view, this width is also the GDD tolerance for the photon pairs to keep coherence, which reflects their time correlation of the biphotons at $\Delta t\!=\!1/\Delta\nu\!\approx\!9$ fs. Together with the measured linewidth of $20$Hz for the SFG peak, mentioned above we obtain $\Delta(\nu_{1}\!+\!\nu_{2})\Delta(t_{1}\!-\!t_{2})\!\approx\!2\!\cdot\!10^{-13}$, showing a giant violation of the classical bound (Eq. 1). The longer-range oscillations in the GDD-scans beyond the central coherent SFG peak are mainly due to the fact that we scan the GDD, rather than the group delay, which indicates that the spectral phase at the center of the spectrum varies quadratically and not linearly across the spectrum.